OpenStudy (cookiemonster18):

PLEASE HELP ME WITH 3 QUESTIONS, I'LL GIVE METALS AND FAN YOU IF YOUR CORRECT 1. starting from 105 feet away a person on a bicycle rides toward a checkpoint and then passes it.the rider is traveling at a constant rate of 35 feet per second.the distance between the bicycle and the checkpoint is given by the equation d=|105-35t| at what time is the bike 60 feet away from the checkpoint? A)1.3 sec and 2.6sec B)1.3sec and 4.7 sec C)1.1 sec and 2.6 sec D)4.7 sec and 9.4 sec

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

\[x=\left| a \right|,x=\pm a\]

OpenStudy (anonymous):

\[105-35 t=\pm 60,find~ t\]

OpenStudy (cookiemonster18):

2.Which number is a solution of the inequality ? 12<y(8-y) A)0 B)1 C)2 D)3

OpenStudy (anonymous):

check one by one

OpenStudy (anonymous):

or \[12< y \left( 8-y \right)\] \[y^2-8y+12<0\] \[y^2-6y-2y+12<0,\left( y-6 \right)\left( y-2 \right)<0\] x lies in (2,6)

OpenStudy (anonymous):

so x=?

OpenStudy (cookiemonster18):

i think 0 because it's a less than sign, im not sure. :/

OpenStudy (cookiemonster18):

@surjithayer

OpenStudy (cookiemonster18):

@Rainbow_Factory can any one please help

OpenStudy (cookiemonster18):

Suppose T = {-8, -4, 0, 4, 8, 12, 16, 20,} and K = {-3, -2, -1, 0,1, 2, 3, 4, 5, 6,} What is T K?

OpenStudy (anonymous):

correction it is y not x

OpenStudy (anonymous):

only 3 lies in (2,6)